use the chain rule to find dz/ds and dz/dt. z=arcsin(x-y), x=s^2+t^2, y=2-6st. dz/ds=? dz/dt=?
Answer #1

Soultion given that 2= arcsin (x-4) sin (2-Y), X= 32+ + 2 and Y = 2-65+ うNow 28 put am at 50 d2 at dz. dx + d2 dy dst at dy adz dt 9+ +63 ST-((S++) d2 dt SHOP 24.63)2 (2++65) |-(s2++2_2+68+)2 Are. = Now we ZP pult ds 50 = d2 ds d2 dx dx SÞ + d2 dy dy